OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = –3x + 12 and passes through the point (−1, 6). (2 points) y = x + 7 y = –3x + 3 y = x + 3 y = –3x + 7

OpenStudy (anonymous):

Well I think it's like this.. y=mx+c 6=-1+c c=7 So to have an equal slope it is y=-3x+7 :)

OpenStudy (anonymous):

Since there is a intercept c could be taken too.

OpenStudy (anonymous):

Thanks